Answer the following questions, which relate to the aggregate expenditures model:

a. Given the following:

     Ca = $120,

     Ig = $60,

     Xn = − $10, and

     G = $40,

What is the economy’s equilibrium GDP?     

Instructions: Enter your answer as a whole number.  

   

     Equilibrium GDP = $  .

b. If real GDP in an economy is currently $240, will the economy’s real GDP rise, fall, or stay the same?  (Click to select)  Real GDP will fall  Real GDP will rise  Real GDP will stay the same  .



c. Suppose that full-employment (and full-capacity) output in an economy is $240. If

     

     Ca = $170,

     Ig = $60,

     Xn = − $10, and

     G = $40,
Answer #1

a.

Ca = $120,

     Ig = $60,

     Xn = − $10, and

     G = $40,

GDP=C+I+G+X-M

=120+60+40+(-10)-0

=$210

Equilibrium GDP is $210.

b.

If real GDP in an economy is currently $240,

Since potential GDP is greater than the actual real GDP, so real GDP will increase until it reaches to its potential level.

C.

Suppose that full-employment (and full-capacity) output in an economy is $240. If

     

     Ca = $170,

     Ig = $60,

     Xn = − $10, and

     G = $40,

GDP=C+I+G+X-M

=170+60+40+(-10)-0

=$260

Equilibrium GDP is $260.

Since real GDP in an economy is currently $240,

Since potential GDP is less than the actual real GDP, so real GDP will decrease until it reaches to its potential level.